# Semantic Communication based on Large Language Model for Underwater Image Transmission

### v1

@@ Abstract
Underwater communication is essential for environmental monitoring, marine biology research, and underwater exploration. Traditional underwater communication faces limitations like low bandwidth, high latency, and susceptibility to noise, while semantic communication (SC) offers a promising solution by focusing on the exchange of semantics rather than symbols or bits. However, SC encounters challenges in underwater environments, including information loss and difficulties in accurately identifying and transmitting critical information that aligns with the diverse requirements of underwater applications. To address these challenges, we propose a novel Semantic Communication (SC) framework based on Large Language Models (LLMs). Our framework leverages visual LLMs to perform semantic compression and prioritization of underwater image data according to the query from users. By identifying and encoding key semantic elements within the images, the system selectively transmits high-priority information while applying higher compression rates to less critical regions. On the receiver side, an LLM-based recovery mechanism, along with Global Vision ControlNet and Key Region ControlNet networks, aids in reconstructing the images, thereby enhancing communication efficiency and robustness. Our framework reduces the overall data size to 0.8% of the original. Experimental results demonstrate that our method significantly outperforms existing approaches, ensuring high-quality, semantically accurate image reconstruction.
@@ I-A Motivation and Background
However, existing research in SC faces significant challenges in the underwater environment. Firstly, the delay spread from multipath propagation and Doppler spread from relative motion lead to information loss in underwater communication. Multipath effects, caused by reflections off the sea surface and seabed, and refractions from varying water densities, distort signals. Frequency-selective fading causes different frequencies to attenuate unevenly, while the Doppler effect from relative motion between transmitters (e.g., underwater vehicles) and receivers (e.g., stationary sensors) introduces frequency shifts. This challenge is particularly problematic for semantic communication, as it hinders the accurate extraction and representation of semantic content, making reliable data transmission and the delivery of meaningful information difficult in underwater environments.
Secondly, SC faces the challenge of accurately identifying and transmitting different critical information in alignment with the diverse requirements of underwater applications. The limitations in extracting and representing semantic content complicate the process of aligning communication with specific application needs. Underwater applications, such as environmental monitoring, marine biology, and exploration, require tailored communication strategies that prioritize different types of data. For instance, environmental monitoring may focus on transmitting images of coral reef health, while marine biology might prioritize detecting specific marine species. Non-essential semantic information can interfere with the transmission of important data, potentially leading to misinterpretations or delays that negatively impact underwater operations and decision-making processes.
@@ I-B2 Semantic Communication
Shannon and Weaver’s[25, 26] work reveals three levels of communication: the lowest level is the symbol transmission level, which focuses on transmitting symbols from the transmitter to the receiver; the intermediate level is the semantic level, where semantic information is extracted at the transmitter, transmitted through a semantic channel, and interpreted at the receiver; and the highest level is the effectiveness level, which ensures the communication efficiency of the other two levels. Traditional communication is measured by the correct bit transmission rate. Due to channel noise and interference, the transmitted symbols can be distorted, leading to the reception of garbled information. On the other hand, SC can extract the meaning of data in the semantic domain, filter out irrelevant and unimportant information, and significantly reduce the data volume by further compressing the semantic information while preserving the meaning. Therefore, when bandwidth is limited or the signal-to-noise ratio (SNR) is relatively low, SC systems may still perform well, maintaining robustness and reliability. Thanks to the development of artificial intelligence and communication technologies, related research has developed end-to-end SC systems based on deep learning to achieve the aforementioned second level of communication. Xie et al. proposed a Transformer-based SC framework to maximize system capacity and eliminate semantic errors[9]. Compared to text, image data is semantically richer and more bandwidth-sensitive, making the transmission of image data more challenging. Huang et al. proposed an image semantic encoding method based on Generative Adversarial Networks (GANs)[27]. Wu et al. proposed a semantic transmission system for segmenting regions of interest in images[28]. Eirina et al. proposed an image transmission method based on joint source-channel coding[29]. However, these methods cannot identify the different critical information that aligns with the diverse requirements of underwater applications, which can lead to the loss of critical information and reduced transmission efficiency. Additionally, multipath propagation and Doppler spread from relative motion in underwater environments further intensify the difficulties in maintaining robust data transmission.
Therefore, it is imperative to adequately consider identifying the varying important information of different tasks when designing SC systems. Moreover, we need to further consider multipath propagation and Doppler spread from relative motion, ensuring that critical information is preserved and accurately reconstructed even under challenging underwater conditions.
@@ I-C Our Contributions
To address these challenges, we propose a novel SC framework based on Large Language Models (LLMs). The advent of LLMs, particularly those designed for visual data, has opened new avenues for enhancing communication systems. Our framework leverages visual LLMs [30] to comprehend queries from individuals above the surface and perform semantic compression and prioritization of image data, aligning with the diverse requirements of underwater applications. This method allows for the identification and encoding of key semantic elements within images and selectively transmits high-priority information while applying higher compression rates to less critical regions, enhancing the effectiveness and relevance of data transmission. On the receiver side, we employ a text LLM recovery mechanism to reconstruct the textual data, and two new ControlNet networks to assist the diffusion model in recovering the image to the original size, thereby mitigating information loss, enhancing the system’s ability to maintain the integrity and detail of the transmitted data. This approach not only reduces the overall data size to 0.8% of the original, but also significantly improves the resilience of the communication system against noise and signal loss.
@@
  • The proposed framework integrates two newly designed ControlNet networks and a text-based LLM recovery mechanism with a diffusion model, ensuring the preservation of key semantic elements while effectively mitigating information loss.
@@ II-B Underwater Acoustic Communication Channel
In this section, we simulate a realistic physical communication environment. In a shallow water acoustic propagation environment, to consider the multipath propagation and Doppler spread, we first define the path loss between surface and underwater robot, which can be defined as:
@@
$\text{SNR}=\frac{P_{signal}}{P_{noise}},$ (8)
@@ III-A Dataset Description and Evaluation Metrics
$d_{F}^{2}(X,Y)=\|\mu_{X}-\mu_{Y}\|_{2}^{2}+Tr(\Sigma_{X}+\Sigma_{Y}-2(\Sigma_{X}\Sigma_{Y})^{\frac{1}{2}}),$ (21)
@@
$SSIM(x,y)=[l(x,y)]^{\alpha}[c(x,y)]^{\beta}[s(x,y)]^{\gamma},$ (22)

### v2

@@ Abstract
Underwater communication is essential for environmental monitoring, marine biology research, and underwater exploration. Traditional underwater communication faces limitations like low bandwidth, high latency, and susceptibility to noise, while semantic communication (SC) offers a promising solution by focusing on the exchange of semantics rather than symbols or bits. However, SC encounters challenges in underwater environments, including semantic information mismatch and difficulties in accurately identifying and transmitting critical information that aligns with the diverse requirements of underwater applications. To address these challenges, we propose a novel Semantic Communication (SC) framework based on Large Language Models (LLMs). Our framework leverages visual LLMs to perform semantic compression and prioritization of underwater image data according to the query from users. By identifying and encoding key semantic elements within the images, the system selectively transmits high-priority information while applying higher compression rates to less critical regions. On the receiver side, an LLM-based recovery mechanism, along with Global Vision ControlNet and Key Region ControlNet networks, aids in reconstructing the images, thereby enhancing communication efficiency and robustness. Our framework reduces the overall data size to 0.8% of the original. Experimental results demonstrate that our method significantly outperforms existing approaches, ensuring high-quality, semantically accurate image reconstruction.
@@ I-A Motivation and Background
However, traditional SC faces significant challenges in underwater environments, particularly for semantic information tasks. Firstly, the delay spread from multipath propagation and Doppler shifts from relative motion leads to semantic information mismatch. Multipath effects and frequency-selective fading distort signals, making it difficult for traditional SC to accurately extract and transmit semantic content. The Doppler effect further exacerbates this by introducing frequency shifts that disrupt the coherence of semantic data. Traditional methods, which rely heavily on a singular recovery mechanism without semantic coherence enhancement, are insufficient to accurately reconstruct meaningful data amidst the complex noise and severe interference of underwater environments.
Secondly, SC faces the challenge of accurately identifying and transmitting different critical information in alignment with the diverse requirements of underwater applications. The traditional SC framework typically lacks the flexibility needed to distinguish and prioritize varying types of data within communication channels. As a result, they struggle to effectively manage the complexity of underwater applications, where different semantic content must be accurately identified, prioritized, and transmitted according to the specific requirements of the task at hand. For instance, environmental monitoring may focus on transmitting images of coral reef health, while marine biology might prioritize detecting specific marine species. Non-essential semantic information can interfere with the transmission of important data, potentially leading to misinterpretations or delays that negatively impact underwater operations and decision-making processes.
@@ I-B2 Semantic Communication
Thanks to the development of artificial intelligence and communication technologies, researchers have developed SC systems based on deep learning to help address the limited bandwidth and robustness of communication. SC can extract the meaning of data in the semantic domain, filter out irrelevant and unimportant information, and significantly reduce the data volume by further compressing the semantic information while preserving the meaning. Therefore, when bandwidth is limited or the signal-to-noise ratio (SNR) is relatively low, SC systems may still perform well, maintaining robustness and reliability. Xie et al. proposed a Transformer-based SC framework to maximize system capacity and eliminate semantic errors[9] while Wang et al. propose the LLM-SC framework that integrates Large Language Models directly into the physical layer of semantic communication systems for text data only[25]. However, compared to text, image data is semantically richer and more bandwidth-sensitive, making the transmission of image data more challenging. Huang et al. proposed an image semantic encoding method based on Generative Adversarial Networks (GANs)[26]. Wu et al. proposed a semantic transmission system for segmenting regions of interest in images[27]. Eirina et al. proposed an image transmission method based on joint source-channel coding[28]. Cao et al. proposed an underwater Metaverse framework utilizing quantum imaging and generative AI to create highly immersive and realistic virtual underwater environments [29] while Jiang et al. propose a LAM-based SC framework for image data that uses AI models to segment and encode semantic data automatically [11]. Zhang et al. also propose a generative AI-aided SC framework that integrates deep generative models for semantic feature extraction and reconstruction [30]. However, these methods cannot identify the different critical information that aligns with the diverse requirements of underwater applications, which can lead to the loss of critical information and reduced transmission efficiency. Additionally, multipath propagation and Doppler spread from relative motion in underwater environments further intensify the difficulties in maintaining robust data transmission.
Therefore, it is imperative to adequately consider identifying the varying important information of different tasks when designing SC systems. Moreover, we need to further consider the semantic information mismatch caused by the complex underwater environments, ensuring that critical information is preserved and accurately reconstructed even under challenging underwater conditions.
@@ I-C Our Contributions
To address these challenges, we propose a novel SC framework based on Large Language Models (LLMs). Unlike traditional models, LLMs excel in their generalization capability, allowing them to be readily applicable across a broad range of tasks within only one model. Beyond this, they can also interpret user intentions to distinguish and prioritize different types of information based on their importance. This enables LLMs to selectively filter and manage critical data according to the specific needs of the task at hand, offering a significant advantage over traditional models. Our framework leverages visual LLMs [31] to comprehend queries from individuals above the surface and perform semantic compression and prioritization of image data, aligning with the diverse requirements of underwater applications. This method allows for the identification and encoding of key semantic elements within images and selectively transmits high-priority information while applying higher compression rates to less critical regions, enhancing the effectiveness and relevance of data transmission. On the receiver side, we employ a text LLM recovery mechanism to reconstruct the textual data, and two new ControlNet networks to assist the diffusion model in recovering the image to the original size. Those recovery mechanisms enhance the semantic coherence to mitigate the information mismatch. This approach not only reduces the overall data size to 0.8% of the original, but also significantly improves the resilience of the communication system against noise and signal loss.
@@
  • The proposed framework integrates two newly designed ControlNet networks and a text-based LLM recovery mechanism with a diffusion model, enhancing the semantic coherence while effectively mitigating information mismatch.
@@ II-B Underwater Acoustic Communication Channel
In this section, a realistic physical communication environment is considered. In a shallow water acoustic propagation environment, to consider the multipath propagation and Doppler spread, we first define the path loss between surface and underwater robot, which can be defined as:
@@
$\text{SNR}=\frac{P_{\text{signal}}}{P_{\text{noise}}},$ (8)
@@ III-A Dataset Description and Evaluation Metrics
$d_{F}^{2}(X,Y)=\|\mu_{X}-\mu_{Y}\|_{2}^{2}+Tr\left[\Sigma_{X}+\Sigma_{Y}-2(\Sigma_{X}\Sigma_{Y})^{\frac{1}{2}}\right],$ (21)
@@
$\text{SSIM}(x,y)=[l(x,y)]^{\alpha}[c(x,y)]^{\beta}[s(x,y)]^{\gamma},$ (22)